# Thresholded Lasso for high dimensional variable selection

### v2

@@ Appendix C Proof sketch of Theorem 3 in [43]
[…]
@@
[…]
@@
However, the largest signal in $H_{0}$ can be $\asymp\sqrt{q}\sigma\lambda$ via the $\eta_{2}$ condition (82):
@@
Then one obtains (86) by (80), (106) and (108). We will compare the conditions and the bounds with Theorem 2.4 in Section 4.3.
C.1 The upper bound in (78) is essentially tight
First, the bound in (77) is tight, since
$\geq$ $(s_{0}-a_{0})(1-C^{2}_{m})\lambda^{2}\sigma^{2}-\lambda^{2}\sigma^{2}.$
$\sum_{j\in A_{0}^{c}}\left\lvert\beta_{j}\right\rvert$ $=$ $\left\lVert\beta^{(2)}\right\rVert_{1}+C_{m}\lambda\sigma(s_{0}-a_{0})=\Omega(\lambda\sigma(s_{0}-a_{0})\frac{\sqrt{(\log p)}}{c_{t}}).$
It remains to show (114), which follows from (74) and
$\left\lVert\beta^{(12)}\right\rVert_{2}^{2}$ $=$ $\sum_{a_{0}<j\leq s_{0}}\min(\beta_{j}^{2},\lambda^{2}\sigma^{2})=C^{2}_{m}\lambda^{2}\sigma^{2}(s_{0}-a_{0}).$
Notation
Recall we use $\beta_{T}\in\mathbb{R}^{\left\lvert T\right\rvert}$, where $T\subseteq[p]$ to also represent its 0-extended version $\beta^{\prime}\in\mathbb{R}^{p}$ such that $\beta^{\prime}_{T^{c}}=0$ and $\beta^{\prime}_{T}=\beta_{T}$. Given $\hat{\beta}_{I}=(X_{I}^{T}X_{I})^{-1}X_{I}^{T}Y\in\mathbb{R}^{\left\lvert I\right\rvert}$, we also use $\hat{\beta}_{I}$ to represent its 0-extended version $\hat{\beta}^{\mathop{\text{\rm ols}\kern 0.60275pt}}\in\mathbb{R}^{p}$ such that $\hat{\beta}^{\mathop{\text{\rm ols}\kern 0.60275pt}}_{I}=\hat{\beta}_{I}$ and $\hat{\beta}^{\mathop{\text{\rm ols}\kern 0.60275pt}}_{I^{c}}=\hat{\beta}_{I^{c}}=0$.
Appendix D Proof of the MSE lower bound
We show the proof of (10) for self-containment. Note that due to different normalization of columns of $X$, our expressions are slightly different from those by [9]. Hence we give a complete derivation here. Consider the least square estimator $\hat{\beta}_{I}=(X_{I}^{T}X_{I})^{-1}X_{I}^{T}Y$, where $|I|\leq s$ and $\hat{\beta}_{I^{c}}=0$, and consider the ideal least-squares estimator $\beta^{\diamond}$ which minimizes the expected mean squared error
$\beta^{\diamond}=\operatorname{argmin}_{I\subset\{1,\ldots,p\},\;|I|\leq s}{\textbf{E}}\left\lVert\beta-\hat{\beta}_{I}\right\rVert_{2}^{2}.$ (116)
Proposition D.1.
@@ Appendix D Proof of the MSE lower bound
Let $I$ be a fixed subset of indices and consider the OLS estimator $\hat{\beta}^{\mathop{\text{\rm ols}\kern 0.60275pt}}$ such that $\hat{\beta}^{\mathop{\text{\rm ols}\kern 0.60275pt}}_{I}=\hat{\beta}_{I}=(X_{I}^{T}X_{I})^{-1}X_{I}^{T}Y$ and $\hat{\beta}^{\mathop{\text{\rm ols}\kern 0.60275pt}}_{I^{c}}=\hat{\beta}_{I^{c}}=0$. Here we again denote by $\beta_{I}$ the restriction of $\beta$ to the set $I$. The error of this estimator is given by
$\left\lVert\hat{\beta}^{\mathop{\text{\rm ols}\kern 0.60275pt}}(I)-\beta\right\rVert_{2}^{2}$ $=$ $\left\lVert\hat{\beta}_{I}-\beta_{I}\right\rVert_{2}^{2}+\left\lVert\beta_{I}-\beta\right\rVert_{2}^{2}.$ (118)
@@
$\hat{\beta}_{I}-\beta_{I}$ ::=:= $(X_{I}^{T}X_{I})^{-1}X_{I}^{T}Y-\beta_{I}$
@@
Appendix E Proof of Lemma 2.8
@@
Appendix F Proof of Lemma 3.1
@@
Appendix G Proof of Lemma 3.3
Thus it follows from the proof Lemma 3.1 that
$\left\lVert h_{T_{01}}\right\rVert_{2}$$\leq$ $\left\lVert Xh_{T_{01}}\right\rVert_{2}/\sqrt{n\Lambda_{\min}(2s_{0})}$
where we replace $\frac{\theta_{s_{0},2s_{0}}}{\sqrt{\Lambda_{\min}(2s_{0})}}$ with $\sqrt{\Lambda_{\max}(s_{0})}$.
Thus we have
$\left\lVert h_{T_{01}}\right\rVert_{2}$$\leq$ $\frac{1}{\sqrt{\Lambda_{\min}(2s_{0})}}\left(\left\lVert Xh\right\rVert_{n}+\frac{\theta_{s_{0},2s_{0}}}{\sqrt{\Lambda_{\min}(2s_{0})}}\left\lVert h_{T_{0}^{c}}\right\rVert_{1}/\sqrt{s_{0}}\right)$
$\left\lVert h_{T_{01}}\right\rVert_{2}$$\leq$ $\frac{1}{\sqrt{\Lambda_{\min}(2s_{0})}}\left(\left\lVert Xh\right\rVert_{n}+\sqrt{\Lambda_{\max}(s_{0})}\left\lVert h_{T_{0}^{c}}\right\rVert_{1}/\sqrt{s_{0}}\right)$
Appendix H Proof of Theorem 2.4
We first show Lemma H.1, which gives us the prediction error using $\beta_{T_{0}}$. We do not focus on obtaining the best constants in the proof of Theorem 2.4. Recall we define a quantity $\lambda_{\sigma,a,p}$, which bounds the maximum correlation between the noise and covariates of $X$; For each $a\geq 0$, let
@@
$\left\lVert X\beta-X\beta_{T_{0}}\right\rVert_{2}/\sqrt{n}$$\leq$ $\sqrt{\Lambda_{\max}(s-s_{0})}\lambda\sigma\sqrt{s_{0}},$ (122)
where recall $\beta_{T_{0}}$ is the the restriction of $\beta$ to the set $T_{0}$.
@@
Hence $\left\lVert X\beta-X\beta_{T_{0}}\right\rVert_{2}/\sqrt{n}=\left\lVert X\beta_{T_{0}^{c}}\right\rVert_{2}/\sqrt{n}\leq\sqrt{\Lambda_{\max}(s-s_{0})}\left\lVert\beta_{T_{0}^{c}}\right\rVert_{2}.$ ∎∎
H.1 Proof of Theorem 2.4: Modern
Prelude. First recall the following elementary inequality. By the optimality of $\hat{\beta}$, we have
$\;\;\;\;\;\frac{1}{2n}\left\lVert Y-X\hat{\beta}\right\rVert^{2}_{2}-\frac{1}{2n}\left\lVert Y-X\beta_{T_{0}}\right\rVert^{2}_{2}\leq\lambda_{n}\left\lVert\beta_{T_{0}}\right\rVert_{1}-\lambda_{n}\left\lVert\hat{\beta}\right\rVert_{1},$ (123)
@@
$=$ $\left\lVert X\hat{\beta}-X\beta\right\rVert_{2}^{2}+2(\beta-\hat{\beta})^{T}X^{T}\epsilon+\left\lVert\epsilon\right\rVert^{2}_{2},$
and similarly, we have for $\beta_{0}=\beta_{T_{0}}$,
@@
This is the inequality we used in [48].
In the updated proof, we will apply the following Lemma H.2; cf. Lemma A.2 [2].
Lemma H.2.
@@
$\leq\left\lVert f-X\tilde{\beta}\right\rVert^{2}_{n}+\frac{2}{n}\xi^{T}X(\hat{\beta}-\tilde{\beta})+h(\tilde{\beta})-h(\hat{\beta})$
Let $\beta_{T_{0}}$ be the the restriction of $\beta$ to the set $T_{0}$. Denote by $h=\hat{\beta}-\beta_{0}$, where $\beta_{0}=\beta_{T_{0}}$. We use $\hat{\beta}:=\beta_{\text{\rm init}}$ to represent the solution to the Lasso estimator in (2). Using Lemma H.2 with $y=f+\epsilon$, where $f=X\beta$, and $\beta_{0}:=\beta_{T_{0}}$, we obtain eq. (20) [11], where we set $\bar{\beta}=\beta_{0}:=\beta_{T_{0}}$ and $\bar{\delta}:=\hat{\beta}-\bar{\beta}=\hat{\beta}-\beta_{0}=h$,
@@
Thus, we have the following updated inequality (40), replacing (H.1). Throughout this proof, we assume that ${\mathcal{T}}_{a}$ holds. Denote by $\delta:=\hat{\beta}-\beta$ and $h=\hat{\beta}-\beta_{0}$. Thus we have by (126), on ${\mathcal{T}}_{a}$,
$\left\lVert X\delta\right\rVert^{2}_{n}+\left\lVert Xh\right\rVert^{2}_{n}$$\leq$ $\left\lVert X\beta-X\beta_{0}\right\rVert^{2}_{n}+\frac{2h^{T}X^{T}\epsilon}{n}+2\lambda_{n}(\left\lVert\beta_{0}\right\rVert_{1}-\left\lVert h+\beta_{0}\right\rVert_{1})$
@@
$\leq$ $\left\lVert X\beta-X\beta_{0}\right\rVert^{2}_{n}+3\lambda_{n}\left\lVert h_{T_{0}}\right\rVert_{1}-\lambda_{n}\left\lVert h_{T_{0}^{c}}\right\rVert_{1},$ (127)
where we have used the fact that $\lambda_{n}\geq 2\lambda_{\sigma,a,p}$ for $a\geq 0$. (40) then follows immediately from (127). We consider three cases which does not need to be mutually exclusive.
@@
$\left\lVert X\delta\right\rVert^{2}_{n}+\left\lVert Xh\right\rVert^{2}_{n}\geq\left\lVert X\beta-X\beta_{0}\right\rVert_{n}^{2}=\left\lVert X\beta_{T_{0}^{c}}\right\rVert^{2}_{n}.$
@@
(45) follows immediately from (129), upon deleting the term $\left\lVert Xh\right\rVert^{2}_{n}$ from both sides. Moreover, we obtain the upper bound on $\left\lVert h_{T_{0}^{c}}\right\rVert_{1}$ ( $\left\lVert h_{T_{0}}\right\rVert$) from (45) since
@@
following the proof of Lemma H.1, where $\lambda_{n}=d_{0}\lambda\sigma\geq 2\lambda_{\sigma,a,p}$. Moreover, we will show that by (129) and (45),
@@
[…]
Then, we have by (128) and (H.1),
@@
$\;\;D_{1,a}$ ::=:= $\big{(}\Lambda_{\max}(s-s_{0})/d_{0}+9d_{0}K(s_{0},3)^{2}/{4}\big{)}\wedge 3D^{\prime}_{0}\;\;\;\text{ where}$
@@
Upper bound on $\left\lVert h_{T_{0}}\right\rVert_{2}$ (44). Under the $\textnormal{{RE}}(s_{0},3,X)$ condition, we have by Definition (3) and (133), for $h=\hat{\beta}-\beta_{0}$, where $\beta_{0}=\beta_{T_{0}}$,
@@
$\left\lvert\left\lVert h_{T_{0}}\right\rVert_{2}-\frac{3K^{2}}{2}\lambda_{n}\sqrt{s_{0}}\right\rvert$$\leq$ $K\left\lVert X\beta-X\beta_{0}\right\rVert_{n}+\frac{3K^{2}}{2}\lambda_{n}\sqrt{s_{0}},\text{ which implies that}$
@@
$D^{\prime}_{0,a}$ ::=:= $K(s_{0},3)\sqrt{\Lambda_{\max}(s-s_{0})}+3K(s_{0},3)^{2}d_{0}.$ (135)
@@
$\left\lVert h\right\rVert_{1}$ $=$ $\left\lVert h_{T_{0}^{c}}\right\rVert_{1}+\left\lVert h_{T_{0}}\right\rVert_{1}\leq\left\lVert X\beta-X\beta_{0}\right\rVert_{n}^{2}/\lambda_{n}+4K^{2}\lambda_{n}s_{0}$
$\leq$ $D_{2,a}\lambda\sigma s_{0},\;\text{ where }\;D_{2,a}\leq\Lambda_{\max}(s-s_{0})/d_{0}+4K(s_{0},3)^{2}d_{0}.$ (136)
@@
where $D_{1,b}=4d_{0}K(s_{0},4)^{2}$ and $D_{2,b}=5d_{0}K(s_{0},4)^{2}$.
Case 3. Suppose
@@
Combining the preceding bound with (144) for Case 2 and (135) for Case 1, we have the expression for (2.1) and the following $\ell_{2}$ error bound on ${h_{T_{0}}}$,
@@
Similarly, combining (136), (146) for Case 2, and (H.1) for Case 3, we obtain the expression of $D_{2}$ in (27):
@@
Proof of Lemma 3.4. Under the settings of Theorem 2.4, we have on ${\mathcal{T}}_{a}$
$\left\lVert X\delta\right\rVert^{2}_{n}+\left\lVert Xh\right\rVert^{2}_{n}+\lambda_{n}\left\lVert h_{T_{0}^{c}}\right\rVert_{1}\leq\left\lVert X\beta_{T_{0}^{c}}\right\rVert^{2}_{n}+3\lambda_{n}\left\lVert h_{T_{0}}\right\rVert_{1}.$ (151)
The rest of the proof is devoted to the $\ell_{2}$ error bound on $h_{T_{01}}$ in view of Lemma 3.1. Let $T_{1}$ be the $s_{0}$ largest positions of $h$ outside of $T_{0}$. Denote by $J_{0}$ the locations of the $s_{0}$ largest coefficients of $h$ in absolute values. Then $J_{0}\subset T_{01}$. Case 1 Now by Proposition A.1 as derived by [46] (151), (44), and Lemma H.1, we have for $K=K(s_{0},3)$,
@@
Then, we have for $h=\hat{\beta}-\beta_{0}$, where $\beta_{0}=\beta_{T_{0}}$, by Lemma 3.1,
@@
[…]
@@
$D$ ::=:= $\frac{\sqrt{\Lambda_{\max}(s-s_{0})}}{\sqrt{\Lambda_{\min}(2s_{0})}}\left(1+\frac{3\ell(s_{0})\sqrt{\Lambda_{\max}(s-s_{0})}}{d_{0}}\right).$ (152)
By Lemmas 3.1,  3.3H.1, Assumption (147), and (52), we obtain for $h=\hat{\beta}-\beta_{0}$,
@@
$\left\lVert h_{T_{0}^{c}}\right\rVert_{1}/\sqrt{s_{0}}$$\leq$ $3\left\lVert X\beta-X\beta_{0}\right\rVert_{n}^{2}/(\lambda_{n}\sqrt{s_{0}})\leq(D_{1,c}\lambda\sigma\sqrt{s_{0}},$
@@
[…]
Combining all three cases, we obtain the expression for (2.1):
@@
It remains to bound $\left\lVert h\right\rVert_{2}$ for Case 3.
Thus we have on ${\mathcal{T}}_{a}$ for $\lambda_{n}=d_{0}\lambda\sigma$ and $\ell(s_{0})$ as defined in (54), for $D$ as in (152), by Lemma 3.1,
@@
Finally, for $h=\hat{\beta}-\beta_{0}$, where $\beta_{0}=\beta_{T_{0}}$, $D_{0}$ as defined in (2.1), and $D_{1}$ as defined in (26), we have by Lemma 3.1,
$\left\lVert\hat{\beta}-\beta\right\rVert_{2}$$\leq$ $\left\lVert\hat{\beta}-\beta_{T_{0}}\right\rVert_{2}+\left\lVert\beta-\beta_{T_{0}}\right\rVert_{2}\leq\left\lVert h\right\rVert_{2}+\lambda\sigma\sqrt{s_{0}}$

### v3

@@ 7 Proof sketch of Theorem 3 in [43]
$(A_{1})\;\;\left|\mathop{\text{\rm supp}\kern 0.86108pt}(\beta_{\text{\rm init}})\cup H_{q}\right|$ $=$ $\left|A_{1}\right|\leq M_{1}^{*}q<q^{*};\text{cf. (2.21) and (3.1)},$ (106)
@@
$\left\lVert X(\beta_{\text{\rm init}}-\beta)\right\rVert_{2}/\sqrt{n}$ $\leq$ $O(1)\sigma\lambda\sqrt{\left\lvert A_{1}\right\rvert}\quad\text{cf. (3.5)}.$ (108)
@@
Then one obtains (86) by (80), (107) and (109). However, the largest signal in $H_{0}$ can be $\asymp\sqrt{q}\sigma\lambda$ via the $\eta_{2}$ condition (82):
@@
We compare the conditions and error bounds with those in Theorem 2.4 in Section 4.3.
8 Conclusion
Acknowledgement
I dedicate this work to Professor Peter Bühlmann, on the occasion of his 60th birthday. His unwavering support and mentorship during my career has been deeply appreciated. The author thanks the Editor, an AE and two anonymous referees for their valuable comments and suggestions. I thank Professor Bin Yu warmly for hosting my visit at UC Berkeley while I finished the first draft of this work in Spring 2010. The author thanks Peter Bühlmann, Emmanuel Candès, Sara van de Geer, John Lafferty, Po-Ling Loh, Richard Samworth, Xiaotong Shen, Martin Wainwright, Larry Wasserman, Dana Yang and Cun-Hui Zhang for helpful discussions. Part of this work has appeared in a conference paper with title: Thresholding Procedures for High Dimensional Variable Selection and Statistical Estimation, in Proceedings of Advances in Neural Information Processing Systems 22.
Appendix A Proof of the MSE lower bound
We show the proof of (10) for self-containment. Note that due to different normalization of columns of $X$, our expressions are slightly different from those by [9]. Hence we give a complete derivation here. Let $I$ be a fixed subset of indices and consider the OLS estimator $\hat{\beta}^{\mathop{\text{\rm ols}\kern 0.60275pt}}$ such that $\hat{\beta}^{\mathop{\text{\rm ols}\kern 0.60275pt}}_{I}=\hat{\beta}_{I}=(X_{I}^{T}X_{I})^{-1}X_{I}^{T}Y$ and $\hat{\beta}^{\mathop{\text{\rm ols}\kern 0.60275pt}}_{I^{c}}=\hat{\beta}_{I^{c}}=0$. Consider the least square estimator $\hat{\beta}_{I}=(X_{I}^{T}X_{I})^{-1}X_{I}^{T}Y$, where $|I|\leq s$ (and $\hat{\beta}_{I^{c}}=0$), and consider the ideal least-squares estimator $\beta^{\diamond}$ which minimizes the expected mean squared error
$\beta^{\diamond}=\operatorname{argmin}_{I\subset\{1,\ldots,p\},\;|I|\leq s}{\textbf{E}}\left\lVert\beta-\hat{\beta}^{\mathop{\text{\rm ols}\kern 0.60275pt}}(I)\right\rVert_{2}^{2}.$ (114)
Proposition A.1.
@@ Appendix A Proof of the MSE lower bound
Here we denote by $\beta_{I}$ the restriction of $\beta$ to the set $I$, where $I\subset[p]$, and $\beta^{\mathop{\natural}\kern 0.60275pt}({I})$ its 0-extended version such that $\beta^{\mathop{\natural}\kern 0.60275pt}_{I^{c}}(I)=0$ and $\beta^{\mathop{\natural}\kern 0.60275pt}_{I}(I)=\beta_{I}$. The error of this OLS estimator, namely, $\hat{\beta}^{\mathop{\text{\rm ols}\kern 0.60275pt}}$ is given by
$\left\lVert\hat{\beta}^{\mathop{\text{\rm ols}\kern 0.60275pt}}(I)-\beta\right\rVert_{2}^{2}$ $=$ $\left\lVert\hat{\beta}_{I}-\beta_{I}\right\rVert_{2}^{2}+\left\lVert\beta^{\mathop{\natural}\kern 0.60275pt}({I})-\beta\right\rVert_{2}^{2}.$ (116)
@@
$\hat{\beta}_{I}-\beta_{I}$ $:=$ $(X_{I}^{T}X_{I})^{-1}X_{I}^{T}Y-\beta_{I}$
@@
Appendix B Proof of Lemma 2.8
@@
Appendix C Proof of Lemma 3.1
@@
Appendix D Proof of Theorem 2.4
Let $\beta_{T_{0}}$ be the restriction of $\beta$ to the set $T_{0}$. Let $h=\beta_{\text{\rm init}}-\beta^{\mathop{\natural}\kern 0.60275pt}({T_{0}})$. Throughout this proof, as a shorthand, we write $h=\hat{\beta}-\beta_{0}$, where $\hat{\beta}=\beta_{\text{\rm init}}$ and $\beta_{0}=\beta^{\mathop{\natural}\kern 0.60275pt}({T_{0}})$ as defined in Lemma 2.3. We first show Lemma D.1, which gives us the prediction error using $\beta_{T_{0}}$. We do not focus on obtaining the best constants in the proof of Theorem 2.4. Recall we define a quantity $\lambda_{\sigma,a,p}$, which bounds the maximum correlation between the noise and covariates of $X$; For each $a\geq 0$, let
@@
$\left\lVert X\beta-X\beta_{0}\right\rVert_{2}/\sqrt{n}$ $\leq$ $\sqrt{\Lambda_{\max}(s-s_{0})}\lambda\sigma\sqrt{s_{0}},$ (120)
where recall $\beta_{0}=\beta^{\mathop{\natural}\kern 0.60275pt}(T_{0})$, as defined in Lemma D.1.
@@
Hence $\left\lVert X\beta-X\beta_{0}\right\rVert_{2}/\sqrt{n}\leq\sqrt{\Lambda_{\max}(s-s_{0})}\left\lVert\beta_{T_{0}^{c}}\right\rVert_{2}.$ ∎∎
Next, we state Lemma D.2. We prove Lemma D.2 in Section F.
Lemma D.2.
Suppose all conditions in Theorem 2.4 hold. Let $T_{1}$ be the $s_{0}$ largest positions of $h$ outside of $T_{0}$. Denote by $T_{01}=T_{0}\cup T_{1}$. Then
Then by Lemma 3.1 and (121), we have
$\;\text{where}\;\ell(s_{0})$ $:=$ $\frac{\theta_{s_{0},2s_{0}}}{\sqrt{\Lambda_{\min}(2s_{0})}}\wedge\sqrt{\Lambda_{\max}(s_{0})}.$
D.1 Proof of Theorem 2.4
Throughout this proof, we assume that ${\mathcal{T}}_{a}$ holds.
Prelude. First recall the following elementary inequality. By the optimality of $\hat{\beta}=\beta_{\text{\rm init}}$, we have
$\;\;\;\;\;\frac{1}{2n}\left\lVert Y-X\hat{\beta}\right\rVert^{2}_{2}-\frac{1}{2n}\left\lVert Y-X\beta_{0}\right\rVert^{2}_{2}\leq\lambda_{n}\left\lVert\beta_{T_{0}}\right\rVert_{1}-\lambda_{n}\left\lVert\hat{\beta}\right\rVert_{1},$ (122)
@@
$=$ $\left\lVert X\hat{\beta}-X\beta\right\rVert_{2}^{2}+2(\beta-\hat{\beta})^{T}X^{T}\epsilon+\left\lVert\epsilon\right\rVert^{2}_{2},\text{ and }$
@@
This is the inequality we used in [48]. In the updated proof, we will apply the following Lemma D.3; cf. Lemma A.2 [2].
Lemma D.3.
@@
[…]
We use $\hat{\beta}:=\beta_{\text{\rm init}}$ to represent the solution to the Lasso estimator in (2). Using Lemma D.3 with $y=f+\epsilon$, where $f=X\beta$, we obtain eq. (20) [11], where we set $\bar{\beta}=\beta_{0}:=\beta^{\mathop{\natural}\kern 0.60275pt}({T_{0}})$ and $\bar{\delta}=\hat{\beta}-\beta_{0}=h$,
@@
Thus, we have the following updated inequality (42), replacing (D.1).
Denote by $\delta:=\hat{\beta}-\beta$ and $h=\hat{\beta}-\beta_{0}$. Thus we have by (125), on ${\mathcal{T}}_{a}$,
$\left\lVert X\delta\right\rVert^{2}_{n}+\left\lVert Xh\right\rVert^{2}_{n}$ $\leq$ $\left\lVert X\beta-X\beta_{0}\right\rVert^{2}_{n}+\frac{2h^{T}X^{T}\epsilon}{n}+2\lambda_{n}(\left\lVert\beta_{0}\right\rVert_{1}-\left\lVert h+\beta_{0}\right\rVert_{1})$ (126)
@@
$\leq$ $\left\lVert X\beta-X\beta_{0}\right\rVert^{2}_{n}+3\lambda_{n}\left\lVert h_{T_{0}}\right\rVert_{1}-\lambda_{n}\left\lVert h_{T_{0}^{c}}\right\rVert_{1},$
where we have used the fact that $\lambda_{n}\geq 2\lambda_{\sigma,a,p}$ for $a\geq 0$. (42) then follows immediately from (126). We consider three cases which do not need to be mutually exclusive.
@@
$\left\lVert X\delta\right\rVert^{2}_{n}+\left\lVert Xh\right\rVert^{2}_{n}\geq\left\lVert X\beta-X\beta_{0}\right\rVert_{n}^{2}.$
@@
Hence (45) follows immediately from (128), upon deleting the term $\left\lVert Xh\right\rVert^{2}_{n}$ from both sides. Moreover, we obtain the upper bound on $\left\lVert h_{T_{0}^{c}}\right\rVert_{1}$ ( $\left\lVert h_{T_{0}}\right\rVert$) from (45) since
@@
following the proof of Lemma D.1, where $\lambda_{n}=d_{0}\lambda\sigma\geq 2\lambda_{\sigma,a,p}$. Moreover, we will show that by (128) and (45),
@@
$\text{ where }\;D^{\prime}_{0}:=(K\sqrt{\Lambda_{\max}(s-s_{0})}+3d_{0}K^{2}).$
Then, we have by (127) and (D.1),
@@
$\;\;D_{1,a}$ $:=$ $\big(\Lambda_{\max}(s-s_{0})/d_{0}+9d_{0}K(s_{0},3)^{2}/{4}\big)\wedge 3D^{\prime}_{0}\;\;\;\text{ where}$
@@
Upper bound on $\left\lVert h_{T_{0}}\right\rVert_{2}$ (44). Under the $\textnormal{{RE}}(s_{0},3,X)$ condition, we have by Definition (3) and (132), for $h=\hat{\beta}-\beta_{0}$, where $\beta_{0}=\beta^{\mathop{\natural}\kern 0.60275pt}({T_{0}})$,
@@
$\left\lvert\left\lVert h_{T_{0}}\right\rVert_{2}-\frac{3K^{2}}{2}\lambda_{n}\sqrt{s_{0}}\right\rvert \displaystyle\leq$ $K\left\lVert X\beta-X\beta_{0}\right\rVert_{n}+\frac{3K^{2}}{2}\lambda_{n}\sqrt{s_{0}},\text{ which implies that}$
@@
$D^{\prime}_{0,a}$ $:=$ $K(s_{0},3)\sqrt{\Lambda_{\max}(s-s_{0})}+3K(s_{0},3)^{2}d_{0}.$ (134)
@@
$\left\lVert h\right\rVert_{1}$ $=$ $\left\lVert h_{T_{0}^{c}}\right\rVert_{1}+\left\lVert h_{T_{0}}\right\rVert_{1}\leq\left\lVert X\beta-X\beta_{0}\right\rVert_{n}^{2}/\lambda_{n}+4K^{2}\lambda_{n}s_{0}$ (135)
$\leq$ $D_{2,a}\lambda\sigma s_{0},\;\text{ where }\;D_{2,a}\leq\Lambda_{\max}(s-s_{0})/d_{0}+4K(s_{0},3)^{2}d_{0}.$
@@
where $D_{1,b}=4d_{0}K(s_{0},4)^{2}$ and $D_{2,b}=5d_{0}K(s_{0},4)^{2}$. Case 3. Suppose
@@
Combining the preceding bound with (143) for Case 2 and (134) for Case 1, we have the expression for (26) and the following $\ell_{2}$ error bound on $h_{T_{0}}$},
@@
Similarly, combining (135), (145) for Case 2, and (D.1) for Case 3, we obtain the expression of $D_{2}$ in (29):
@@
Appendix E Proof of Lemma 3.3
Let $h=\beta_{\text{\rm init}}-\beta^{\mathop{\natural}\kern 0.60275pt}({T_{0}})$. Throughout this proof, as a shorthand, we write $h=\hat{\beta}-\beta_{0}$, where $\hat{\beta}=\beta_{\text{\rm init}}$ and $\beta_{0}=\beta^{\mathop{\natural}\kern 0.60275pt}({T_{0}})$ as defined in Lemma 2.3. Under the settings of Theorem 2.4, we have on ${\mathcal{T}}_{a}$,
$\left\lVert X\delta\right\rVert^{2}_{n}+\left\lVert Xh\right\rVert^{2}_{n}+\lambda_{n}\left\lVert h_{T_{0}^{c}}\right\rVert_{1}\leq\left\lVert X(\beta-\beta_{0})\right\rVert^{2}_{n}+3\lambda_{n}\left\lVert h_{T_{0}}\right\rVert_{1}.$ (150)
The rest of the proof is devoted to the $\ell_{2}$ error bound on $h_{T_{01}}$ in view of Lemma 3.1. Let $T_{1}$ be the $s_{0}$ largest positions of $h$ outside of $T_{0}$. Denote by $J_{0}$ the locations of the $s_{0}$ largest coefficients of $h$ in absolute values. Then $J_{0}\subset T_{01}$. Case 1 Now by Proposition A.1 as derived by [46], (150), (44), and Lemma D.1, we have for $K=K(s_{0},3)$,
@@
By Lemma 3.1,
@@
$\text{ and }\;\left\lVert h\right\rVert_{2}^{2}$ $\leq$ $(1+k_{0})\left\lVert h_{T_{01}}\right\rVert_{2}^{2},\;\text{ where}\;h\in{\mathcal{C}}(s_{0},k_{0}),$
@@
$D$ $:=$ $\frac{\sqrt{\Lambda_{\max}(s-s_{0})}}{\sqrt{\Lambda_{\min}(2s_{0})}}\left(1+\frac{3\ell(s_{0})\sqrt{\Lambda_{\max}(s-s_{0})}}{d_{0}}\right).$ (151)
By Lemmas 3.1,  D.2D.1, Assumption (146), and (51), we obtain for $h=\hat{\beta}-\beta_{0}$,
@@
$\left\lVert h_{T_{0}^{c}}\right\rVert_{1}/\sqrt{s_{0}}$ $\leq$ $3\left\lVert X\beta-X\beta_{0}\right\rVert_{n}^{2}/(\lambda_{n}\sqrt{s_{0}})\leq D_{1,c}\lambda\sigma\sqrt{s_{0}},$
@@
$\text{ and }\;\left\lVert Xh\right\rVert_{n}$ $\leq$ $\left\lVert X(\beta-\beta_{0})\right\rVert_{n}\leq\sqrt{\Lambda_{\max}(s-s_{0})}\lambda\sigma\sqrt{s_{0}}.$
Combining all three cases, we obtain the expression for (23):
@@
It remains to bound $\left\lVert h\right\rVert_{2}$ for Case 3. Thus we have on ${\mathcal{T}}_{a}$ for $\lambda_{n}=d_{0}\lambda\sigma$ and $\ell(s_{0})$ as defined in (D.2), for $D$ as in (151), by Lemma 3.1,
@@
Finally, for $h=\hat{\beta}-\beta_{0}$, where $\beta_{0}=\beta^{\mathop{\natural}\kern 0.60275pt}({T_{0}})$, $D_{0}$ as defined in (23), and $D_{1}$ as defined in (27), we have by the triangle inequality and Lemma 3.1,
$\left\lVert\hat{\beta}-\beta\right\rVert_{2}$ $\leq$ $\left\lVert\hat{\beta}-\beta_{0}\right\rVert_{2}+\left\lVert\beta-\beta_{0}\right\rVert_{2}\leq\left\lVert h\right\rVert_{2}+\lambda\sigma\sqrt{s_{0}}$